OpenStudy (anonymous):

help please -6x=10-y write it as a functional notation

OpenStudy (anonymous):

function notation includes f(x) f(x) = mx + b is usually function notation so f(x) would represent y, so solve in terms of y and then replace y with f(x) -6x = 10-y y = 10+6x f(x) = 6x + 10